Pyridine, an organic base, has the formula C5H5N , has a KB = 1.78 x 10–9.
Write the equilibrium formed when C5H5N dissolves in water.
Identify the conjugate acid of C5H5N and calculate its KA.
1
Expert's answer
2020-05-07T14:16:14-0400



C5H5NH+

pKb = -log (1.78 x 10–9) = 8.75

pKa + pKb = 14

pKa = 14 - 8.75 = 5.25

Ka = 5,62x10-6